# Competitive Equilibrium for Chores: from Dual Eisenberg-Gale to a Fast, Greedy, LP-based Algorithm

### v1

@@ Appendix A A Primal-Dual Analysis of the Convergence of the Frank Wolfe Method to an Approximate Equilibrium
Note that $x^{t-1}_{ij}>0$ only if $d_{ij}/p_{j}$ is minimum for all $j\in[m]$ for agent $i$. Thus, if $(p^{t-1},x^{t-1})$ is not an $\varepsilon$-approximate equilibrium, then there exists a $j\in[m]$ such that $\sum_{i\in[n]}x^{t-1}_{ij}<1-\varepsilon$, or there exists an $i\in[n]$, such that $|\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}-B_{i}|>B_{i}\varepsilon$.
  • Case $\sum_{i\in[n]}x^{t-1}_{ij}<1-\varepsilon$: From Claim 3, it follows that $\sum_{i\in[n]}x^{t-1}_{ij}=\frac{1}{B}\cdot\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<1-\varepsilon$, implying that $\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<B(1-\varepsilon)$. Further, observe that
    $\sum_{j\in[n]}p^{t-1}_{j}\cdot x^{t-1}_{ij}$ $=\beta^{t-1}_{i}\cdot\sum_{j\in[m]}d_{ij}x^{t-1}_{ij}$ (by Claim 2)
    $=\beta^{t-1}_{i}\cdot\frac{B_{i}}{\beta^{t-2}_{i}}$ (by Claim 1)
    Thus, we have $\sum_{i\in[n]}\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<B(1-\varepsilon)$. This implies that $\prod_{i\in[n]}(\sum_{j\in[m]}$ $p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}\leq\prod_{i\in[n]}(B_{i}(1-\varepsilon))^{B_{i}}$, further implying that
    $\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$ $\geq\frac{1}{(1-\varepsilon)^{B}}$
    $\geq(1+\varepsilon)^{B}$
    $\geq 1+{B\varepsilon}$
    Therefore, $\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}\geq\log(1+B\varepsilon)\geq\min_{i}B_{i}\varepsilon^{2}/2.3$ as $\log(1+x)\geq x^{2}/2.3$ for all $x\in[0,1]$.
  • Case $|\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}-B_{i}|>B_{i}\varepsilon$: Let $i$ be an agent such that $\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=B_{i}(1+\varepsilon^{\prime})$, where $\varepsilon^{\prime}\in\{-\theta,\theta\}$ for some $\theta>\varepsilon$. Observe that $\sum_{\ell\in[n]\setminus\{i\}}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{\ell j})<B-B_{i}-B_{i}\varepsilon^{\prime}$, as $\sum_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})=\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}\leq B$, and $\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=B_{i}(1+\varepsilon^{\prime})$. This implies that $\prod_{\ell\in[n]\setminus\{i\}}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{\ell j})^{B_{i}}\leq\prod_{\ell\in[n]\setminus\{i\}}(B_{\ell}\cdot(1-\frac{B_{i}\varepsilon^{\prime}}{B-B_{i}}))^{B_{\ell}}$. Before, we proceed, we mention the following useful fact,
    Fact 5 ().
@@
    $\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$ $>\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{(B_{i}(1+\varepsilon^{\prime}))^{B_{i}}\prod_{\ell\in[n]\setminus\{i\}}(B_{\ell}\cdot(1-\frac{B_{i}\varepsilon^{\prime}}{B-B_{i}}))^{B_{\ell}}}$
    $=\frac{1}{(1+\varepsilon^{\prime})^{B_{i}}(1-\frac{B_{i}\varepsilon^{\prime}}{B-B_{i}})^{(B-B_{i})}}$
    $\geq\frac{1}{(1+\varepsilon^{\prime})^{B_{i}}e^{-\frac{B_{i}\varepsilon^{\prime}}{B-B_{i}}\cdot(B-B_{i})}}$
    $=\frac{1}{(1+\varepsilon^{\prime})^{B_{i}}e^{-B_{i}\varepsilon^{\prime}}}$
    $=\bigg{(}\frac{1}{(1+\varepsilon^{\prime})e^{-\varepsilon^{\prime}}}\bigg{)}^{B_{i}}\qed$
@@
$\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}$ $>\log\Bigg{(}\bigg{(}\frac{1}{(1+\varepsilon^{\prime})e^{-\varepsilon^{\prime}}}\bigg{)}^{B_{i}}\Bigg{)}$
$=B_{i}\log\Bigg{(}\frac{e^{\varepsilon^{\prime}}}{1+\varepsilon^{\prime}}\Bigg{)}$
$\geq B_{i}\frac{\varepsilon^{\prime 2}}{2.3}$ (from Fact 5)
$=\min_{i}B_{i}\frac{\theta^{2}}{2.3}=\min_{i}B_{i}\frac{\varepsilon^{2}}{2.3}.$
@@
Theorem 6 ().
Let $T\geq\frac{2.3(f(\beta^{*})-f(\beta^{0}))}{\varepsilon^{2}\min_{i}B_{i}}$. Then there exists a $t\leq T$, such that $(p^{t},x^{t})$ is a $\varepsilon$-CE
@@
Assume otherwise, i.e., for all $t\in[T]$, $(p^{t},x^{t})$ is not a $\varepsilon$-CE. Recall that we have,
@@
By Claim 1, we can replace $\sum_{j\in[m]}d_{ij}x^{t}_{ij}$ with $1/\beta^{t-1}_{i}$, implying,
(21) $\frac{\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}}}{\prod_{i\in[n]}(1/\beta^{t-2}_{i})^{B_{i}}}=\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$
Since $(p^{t-1},x^{t-1})$ is not a $\varepsilon$-CEEI, we have $\log\bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\bigg{)}\geq\min_{i}B_{i}\cdot\varepsilon^{2}/2.3$ by Claim 4. Taking logs on both sides of Equation 21, we have
(22) $f(\beta^{t-1})-f(\beta^{t-2})\geq\min_{i}B_{i}\cdot\varepsilon^{2}/2.3$
Summing up Equation 22 from $t=2$ to $T+1$, we have $f(\beta^{T})-f(\beta^{0})>\min_{i}B_{i}\cdot T\varepsilon^{2}/2.3$. Since $f(\beta^{*})\geq f(\beta^{T})$, this implies that $f(\beta^{*})-f(\beta^{0})>\min_{i}B_{i}\cdot T\varepsilon^{2}/2.3$, implying that $T<2.3(f(\beta^{*})-f(\beta^{0}))/(\varepsilon^{2}\cdot\min_{i}B_{i})$, which is a contradiction. ∎
@@ Appendix B Approximate CE and Exact CE can be far Apart
We show that approximate CE and Exact CE can be far apart. Consider the following chores market: 2 agents $a_{1}$ and $a_{2}$, and 2 chores $c_{1}$ and $c_{2}$. Set $d_{11}=1$, and $d_{12}=M$ for $M\gg 1$. Set $d_{21}=1-\varepsilon$, and $d_{22}=1+\varepsilon$ for $\varepsilon\ll 1$. It is easy to verify that the only CEEI here is
@@
Note that $\beta_{1}=2/(M+1)$, and $\beta_{2}=2M/((M+1)(1+\varepsilon))$. However, the following is an $\varepsilon$-approximate CEEI, with $\beta_{1}=\beta_{2}=1$, which is far from the only exact CEEI!
  • $p_{1}=1-\varepsilon$ and $p_{2}=1+\varepsilon$,
  • $x_{11}=1-\varepsilon$, $x_{21}=0$, and

### v2

@@ Appendix A A Primal-Dual Analysis of the Convergence of the Frank Wolfe Method to an Approximate Equilibrium
Note that $x^{t-1}_{ij}>0$ only if $d_{ij}/p_{j}$ is minimum for all $j\in[m]$ for agent $i$. Thus, if $(p^{t-1},x^{t-1})$ is not an $\epsilon$-approximate equilibrium, then there exists a $j\in[m]$ such that $\sum_{i\in[n]}x^{t-1}_{ij}<1-\epsilon$, or there exists an $i\in[n]$, such that $|\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}-B_{i}|>B_{i}\epsilon$.
  • Case $\sum_{i\in[n]}x^{t-1}_{ij}<1-\epsilon$: From 4, it follows that $\sum_{i\in[n]}x^{t-1}_{ij}=\frac{1}{B}\cdot\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<1-\epsilon$, implying that $\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<B(1-\epsilon)$. Further, observe that
    $\sum_{j\in[n]}p^{t-1}_{j}\cdot x^{t-1}_{ij}$ $=\beta^{t-1}_{i}\cdot\sum_{j\in[m]}d_{ij}x^{t-1}_{ij}$ (by 3)
    $=\beta^{t-1}_{i}\cdot\frac{B_{i}}{\beta^{t-2}_{i}}$ (by 2)
    Thus, we have $\sum_{i\in[n]}\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}<B(1-\epsilon)$. This implies that $\prod_{i\in[n]}(\sum_{j\in[m]}$ $p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}\leq\prod_{i\in[n]}(B_{i}(1-\epsilon))^{B_{i}}$, further implying that
    $\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$ $\geq\frac{1}{(1-\epsilon)^{B}}$
    $\geq(1+\epsilon)^{B}$
    $\geq 1+{B\epsilon}$
    Therefore, $\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}\geq\log(1+B\epsilon)\geq\min_{i}B_{i}\epsilon^{2}/2.3$ as $\log(1+x)\geq x^{2}/2.3$ for all $x\in[0,1]$.
  • Case $|\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}-B_{i}|>B_{i}\epsilon$: Let $i$ be an agent such that $\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=B_{i}(1+\epsilon^{\prime})$, where $\epsilon^{\prime}\in\{-\theta,\theta\}$ for some $\theta>\epsilon$. Observe that $\sum_{\ell\in[n]\setminus\{i\}}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{\ell j})<B-B_{i}-B_{i}\epsilon^{\prime}$, as $\sum_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})=\sum_{i\in[n]}B_{i}\frac{\beta^{t-1}_{i}}{\beta^{t-2}_{i}}\leq B$, and $\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij}=B_{i}(1+\epsilon^{\prime})$. This implies that $\prod_{\ell\in[n]\setminus\{i\}}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{\ell j})^{B_{i}}\leq\prod_{\ell\in[n]\setminus\{i\}}(B_{\ell}\cdot(1-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}}))^{B_{\ell}}$. Before, we proceed, we mention the following useful fact,
    Fact 1.
@@
    $\hskip 20.0pt\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$
    $>\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{(B_{i}(1+\epsilon^{\prime}))^{B_{i}}\prod_{\ell\in[n]\setminus\{i\}}(B_{\ell}\cdot(1-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}}))^{B_{\ell}}}$
    $=\frac{1}{(1+\epsilon^{\prime})^{B_{i}}(1-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}})^{(B-B_{i})}}$
    $\geq\frac{1}{(1+\epsilon^{\prime})^{B_{i}}e^{-\frac{B_{i}\epsilon^{\prime}}{B-B_{i}}\cdot(B-B_{i})}}$
    $=\frac{1}{(1+\epsilon^{\prime})^{B_{i}}e^{-B_{i}\epsilon^{\prime}}}$
    $=\bigg{(}\frac{1}{(1+\epsilon^{\prime})e^{-\epsilon^{\prime}}}\bigg{)}^{B_{i}}$
@@
$\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}$ $>\log\Bigg{(}\bigg{(}\frac{1}{(1+\epsilon^{\prime})e^{-\epsilon^{\prime}}}\bigg{)}^{B_{i}}\Bigg{)}$
$=B_{i}\log\Bigg{(}\frac{e^{\epsilon^{\prime}}}{1+\epsilon^{\prime}}\Bigg{)}$
$\geq B_{i}\frac{\epsilon^{\prime 2}}{2.3}$ (from Fact 1)
$=\min_{i}B_{i}\frac{\theta^{2}}{2.3}=\min_{i}B_{i}\frac{\epsilon^{2}}{2.3}.$
@@
Theorem 8.
Let $T\geq\frac{2.3(f(\beta^{*})-f(\beta^{0}))}{\epsilon^{2}\min_{i}B_{i}}$. Then there exists a $t\leq T$, such that $(p^{t},x^{t})$ is a $\epsilon$-CE
@@
Assume otherwise, i.e., for all $t\in[T]$, $(p^{t},x^{t})$ is not a $\epsilon$-CE. Recall that we have,
@@
By 2, we can replace $\sum_{j\in[m]}d_{ij}x^{t}_{ij}$ with $1/\beta^{t-1}_{i}$, implying,
$\frac{\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}}}{\prod_{i\in[n]}(1/\beta^{t-2}_{i})^{B_{i}}}=\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}$ (20)
Since $(p^{t-1},x^{t-1})$ is not a $\epsilon$-CEEI, we have $\log\bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\bigg{)}\geq\min_{i}B_{i}\cdot\epsilon^{2}/2.3$ by 5. Taking logs on both sides of Eq. 20, we have
$f(\beta^{t-1})-f(\beta^{t-2})\geq\min_{i}B_{i}\cdot\epsilon^{2}/2.3$ (21)
Summing up Eq. 21 from $t=2$ to $T+1$, we have $f(\beta^{T})-f(\beta^{0})>\min_{i}B_{i}\cdot T\epsilon^{2}/2.3$. Since $f(\beta^{*})\geq f(\beta^{T})$, this implies that $f(\beta^{*})-f(\beta^{0})>\min_{i}B_{i}\cdot T\epsilon^{2}/2.3$, implying that $T<2.3(f(\beta^{*})-f(\beta^{0}))/(\epsilon^{2}\cdot\min_{i}B_{i})$, which is a contradiction. ∎
Remark.
This analysis also gives us additional insights: for instance, any iteration where the solution of the GFW is not a good approximation of a CE, will see a large improvement in the objective function, implying that GFW iterations with bad approximate solutions are small.
@@ Appendix B Approximate CE and Exact CE can be far Apart
We show that approximate CE and Exact CE can be far apart. Consider the following chores market: 2 agents $a_{1}$ and $a_{2}$, and 2 chores $c_{1}$ and $c_{2}$. Set $d_{11}=1$, and $d_{12}=M$ for $M\gg 1$. Set $d_{21}=1-\epsilon$, and $d_{22}=1+\epsilon$ for $\epsilon\ll 1$. It is easy to verify that the only CEEI here is
@@
Note that $\beta_{1}=2/(M+1)$, and $\beta_{2}=2M/((M+1)(1+\epsilon))$. However, the following is an $\epsilon$-approximate CEEI, with $\beta_{1}=\beta_{2}=1$, which is far from the only exact CEEI!
  • $p_{1}=1-\epsilon$ and $p_{2}=1+\epsilon$,
  • $x_{11}=1-\epsilon$, $x_{21}=0$, and
@@
Appendix C Additional numerical experiments
Figure 5: Numerical results for 100 buyers and up to 1000 chores. GFW solves every instance for every distribution.
Figure 6: Numerical results with very small Gurobi tolerance parameters.